# SOUS VIDE: Cooking Visual Drone Navigation Policies in a Gaussian Splatting Vacuum

### v1

@@ VII Conclusions
This work introduces the SOUS VIDE approach for training end-to-end visual drone navigation policies. SOUS VIDE comprises the FiGS simulator based on a Gaussian Splatting scene model, data collection from an MPC expert, and distillation into a lightweight visuomotor policy architecture. By coupling high-fidelity visual data synthesis with online adaptation mechanisms, SOUS VIDE achieves zero-shot sim-to-real transfer, demonstrating robustness to variations in mass, thrust, lighting, and dynamic scene changes. Our experiments underscore the policy’s ability to generalize across diverse scenarios, including complex and extended trajectories, with graceful degradation under extreme conditions. Notably, the integration of a streamlined adaptation module enables the policy to overcome limitations of prior visuomotor approaches, offering a computationally efficient yet effective solution for addressing model inaccuracies. These findings highlight the potential of SOUS VIDE as a foundation for future advancements in autonomous drone navigation.
Limitations and Future Work: While its robustness and versatility are evident, challenges such as inconsistent performance in multi-objective tasks suggest opportunities for improvement through more sophisticated objective encodings. SOUS VIDE has been used to train policies that are highly optimized for a single real-life environment. Future work will explore training polices with the same tools across multiple environments in FiGS to enable generalist skills, like general collision avoidance, and scene-agnostic navigation. We will also explore augmenting SOUS VIDE policies with semantic goal understanding, so goals can be given by a human operator in the form of natural language commands. Ultimately, this work paves the way for deploying learned visuomotor policies in real-world applications, bridging the gap between simulation and practical autonomy in drone operations.

### v2

@@ VII Conclusions
This work introduces the SOUS VIDE approach for training end-to-end visual drone navigation policies. SOUS VIDE comprises the FiGS simulator based on a Gaussian Splat scene model, data generation from a simulated MPC expert, and distillation into a lightweight visuomotor policy architecture. By coupling high-fidelity visual data synthesis with online adaptation mechanisms, SOUS VIDE achieves zero-shot sim-to-real transfer, demonstrating robustness to variations in mass, thrust, lighting, and dynamic scene changes. Our experiments underscore the policy’s ability to generalize across diverse scenarios, including complex and extended trajectories, with graceful degradation under extreme conditions. Notably, the integration of a streamlined adaptation module enables the policy to overcome limitations of prior visuomotor approaches, offering a computationally efficient yet effective solution for addressing model inaccuracies. These findings highlight the potential of SOUS VIDE as a foundation for future advancements in autonomous drone navigation.
Limitations and Future Work: While its robustness and versatility are evident, challenges such as inconsistent performance in multi-objective tasks suggest opportunities for improvement through more sophisticated objective encodings. SOUS VIDE has been used to train policies that are highly optimized for a single real-life environment. Future work will explore training policies with the same tools across multiple environments in FiGS to enable generalist skills, like general collision avoidance, and scene-agnostic navigation. We will also explore augmenting SOUS VIDE policies with semantic goal understanding, so goals can be given by a human operator in the form of natural language commands. Ultimately, this work paves the way for deploying learned visuomotor policies in real-world applications, bridging the gap between simulation and practical autonomy in drone operations.